OpenStudy (hesan):

Find the two numbers whose sum is 8, if the product of one number and the cube of the other is a maximum value?

ganeshie8 (ganeshie8):

this is calculus.. .?

OpenStudy (anonymous):

yup yet is the answer is 2,6

OpenStudy (hesan):

HOW?

OpenStudy (hesan):

@matricked: u r right, can u please explain me..?

OpenStudy (anonymous):

if one number is x then u have to maximise x^3(8-x) on simply differentiating u get the values either 0 or 6 and then test for max condition using second order

OpenStudy (hesan):

What is meant by maximum value here?

OpenStudy (anonymous):

it means max value of x^3(8-x)

OpenStudy (hesan):

Okey, how will we maximise x^3(8-x)? By equating it to what?

OpenStudy (anonymous):

if u equate it to something then u willsolve for such particular values of x satisfying ur choice u need to differentiate x^3(8-x) and obtain 24x2-4x3and equate it to0 to obtain the values of x(extremums) as x=0,6

OpenStudy (hesan):

i m not getting it...

OpenStudy (anonymous):

if u equate 24x2-4x3=0 thn the values of x is 0 or 6....

OpenStudy (hesan):

24x2-4x3=0 or 24x^2-4x^3=0 ?

OpenStudy (anonymous):

24x^2-4x^3=0 ?

OpenStudy (hesan):

where are the powers in 24x2-4x3=0

OpenStudy (anonymous):

generally its understood...

OpenStudy (hesan):

if we multiply 3x^2(8-x) then we'll get 24x^2-3x^3..